Solve the folowing system.
x - y = 1
x + y = 1

Answer by funmath(2933)

Solve the folowing system.
x - y = 1
x + y = 1
This is the elimination method, add the two equations and the y is eliminated.
x-y=1
x+y=1
______
2x+0=2
2x=2
2x/2=2/2
x=1
Substitute 1 in for x in either equation and solve for y.
1+y=1
-1+1+y=1-1
y=0
The solution (x,y)=(1,0)
Happy Calculating!!!